OpenStudy (anonymous):

At the city museum, child admission is $5.40 and adult admission is $8.90 . On Thursday, 160 tickets were sold for a total sales of $1182.50 . How many adult tickets were sold that day?

OpenStudy (bloomlocke367):

okay, you're going to have to write a system of equations, do you know how to do that?

OpenStudy (bloomlocke367):

@Cassandra_Lea_96

OpenStudy (anonymous):

right now I have 5.40(160-x) + 8.90x = 1182.50

OpenStudy (anonymous):

I'm pretty sure it's supposed to go along those lines not sure I have it set up in the right order though =o

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (bloomlocke367):

I'm not sure if that works or not, it might, but what I would have done is make a system of equations that's \(5.40x+8.90y=1182.50\) and \(x+y=160\)

OpenStudy (bloomlocke367):

do you understand why that will help you solve your question?

OpenStudy (anonymous):

oh ok yes, because x and y are how many tickets are sold and the total of tickets add up to 160. how would I start to work this out?

OpenStudy (bloomlocke367):

you could do a few things, but the easiest (in my opinion) would be to use substitution

OpenStudy (bloomlocke367):

do you know how to do that?

OpenStudy (bloomlocke367):

@Cassandra_Lea_96

OpenStudy (anonymous):

I'd need to work out the x+y = 160 equation to find x and then substitute it right?

OpenStudy (bloomlocke367):

yes :)

OpenStudy (anonymous):

5.40 (160 - y) + 8.909y = 1182.50

OpenStudy (bloomlocke367):

yes, that's really close to what you had :)

OpenStudy (bloomlocke367):

now just solve it, do you know how?

OpenStudy (anonymous):

864 - 5.40y +8.90y = 1182.50 864 - 14.03y = 1182.50 14.03y = 318.05 y = 22.70

OpenStudy (anonymous):

I think I messed up somewhere or missed a step because 22.70 doesn't look like the answer to how many adult tickets were sold =o

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (bloomlocke367):

I see your error... when you combined like terms, you made a mistake.. you got -14.03y when it should be 3.50y

OpenStudy (bloomlocke367):

do you see why?

OpenStudy (anonymous):

oh I should have kept the 5.40y as a negative

OpenStudy (bloomlocke367):

yes :)

OpenStudy (bloomlocke367):

now try it, and see what you get c:

OpenStudy (anonymous):

90.87 ? =p

OpenStudy (bloomlocke367):

I got 91 ... how'd you get 90.87?

OpenStudy (anonymous):

no idea haha calculator? =o I mean it rounds up though I guess =/

OpenStudy (bloomlocke367):

Yeah XD

OpenStudy (anonymous):

lol thank you!!

OpenStudy (bloomlocke367):

You're welcome!